if you have a 4 degree angle and one side (a) equals 7 what is the other side? (b)
1
Expert's answer
2013-03-11T11:39:05-0400
Since angle a=4 degrees, then the other angle will be equalto b=90-4=86 degrees.

If side A=7, then tan(a)=A/B, where B is the other side.
So, B has to be equalto B=A/tan(4) = 7 / tan(4) = 100.105 .